• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name paschke

    Log entry time 12:41:10 on August28,2009

    Entry number 286570

    keyword=BCM calibration runs, BPM calibration run

    11693 - last run from Luis.  Everything stable, checking levels.
    BCM 1 (upstream 1x), BCM 2 (upstream 1x), BCM 5 (unser) all ready to go
    11694 - BCM calibration around 100 uA.
    11695 - BCM calibration.  Attenuator set points between 132 (~100uA) and 124 
    (~80 uA).  In order, attenuator set points called were:
    124,127,132,129,128,126,132,125,127.  Set to beamsync in between for short 
    pedestal measurements.
    **NOTE: We need to redo the calibration soon, with the counting DAQ, for the 
    PVDIS tests. This cannot be done at the same time unless we split the Unser 
    signal, which right now is only going to the parity DAQ.
    11696 - BPM calibration run.  Varied current, with bpms in "gain off" at 100 
    uA.  I couldn't screen grab the bpm live displays with the gain off settings, 
    so I asked MCC to put those in the elog.